# Let’s Take Esoteric Programming Languages Seriously

### v1

@@ Abstract.
Esoteric programming languages are challenging to learn, but their unusual features and constraints may serve to improve programming ability. From languages designed to be intentionally obtuse (e.g. INTERCAL) to others targeting artistic expression (e.g. Piet) or exploring the nature of computation (e.g. Fractan), there is rich variety in the realm of esoteric programming languages. This essay examines the counterintuitive appeal of esoteric languages and seeks to analyse reasons for this popularity. We will explore why people are attracted to esoteric languages in terms of (a) program comprehension and construction, as well as (b) language design and implementation. Our assertion is that esoteric languages can improve general PL awareness, at the same time as enabling the esoteric programmer to impress their peers with obscure knowledge. We will also consider pedagogic principles and the use of AI, in relation to esoteric languages. Emerging from the specific discussion, we identify a general set of ‘good’ reasons for designing new programming languages. It may not be possible to be exhaustive on this topic, and it is certain we have not achieved that goal here. However we believe our most important contribution is to draw attention to the varied and often implicit motivations involved in programming language design.
conference: ; 2025;
@@ 1. Introduction
Despite their name, esoteric programming languages are surprisingly popular in the general Computer Science community. This essay is based on a presentation we have given several times to audiences of academic Computer Scientists; on each occasion we have encountered keen advocates of esoteric languages like Whitespace and Brainf***. In this work, we seek to analyse potential reasons for the broad appeal of such esoteric languages. We will argue that Computer Scientists need to take esoteric languages seriously.
@@
An esoteric programming language (hereafter contracted to esolang, following general convention) is ‘a programming language designed to experiment with weird ideas, to be hard to program in, or as a joke, rather than for practical use’. [^1] https://esolangs.org/wiki/Esoteric_programming_language Temkin (Temkin, 2017) characterises an esoteric programming language as one that is ‘intentionally ununsable, uncomputable or conceptual’. Another definition [^2] https://esoteric.codes/blog/esolangs-as-an-experiential-practice is that ‘esolangs are a class of languages made for reasons other than practical use’.
@@
The claim of this essay is that esoteric languages are inherently useful, and deserve to be studied in a serious manner. The first question we seek to address is: What is the appeal of esolangs, in general? Our discussion will cover the following four high-level reasons for their popularity:
@@
Taking inspiration from the conceptual framework of Computational Thinking (Wing, 2006), we will ask: what do programmers actually do? We will explore how this characteristic behaviour is manifested in the medium of esolangs. Whereas Wing’s claim is that practitioners in other disciplines can benefit from a grasp of computing principles, our study of esoteric languages leads us to believe that computer programmers can benefit from a wider appreciation of other disciplines. This is Gordon’s argument in his Onward! essay from last year (Gordon, 2024) which takes the opposite stance to Wing.
With consideration to pedagogic principles, we will ask Why is it beneficial to expose learners to esoteric languages? We will explore this question through the lens of variation theory, inter alia.
Although esolangs may seem trivial, we feel they deserve our attention since they may change the way we think about programming and language design. As Perlis (Perlis, 1982) puts it:
(19) A language that doesn’t affect the way you think about programming, is not worth knowing.
@@ 2. Background
In this section, we briefly survey the field of esolangs. We commence with the prototypical esolang, INTERCAL, in Section 2.1. After this, we group the languages into related sets, considering esolangs that have a natural language syntax (Section 2.2), those with exotic syntax (Section 2.3) and those with an exotic computational model (Section 2.4). We acknowledge that these characteristic properties are not mutually exclusive. Further, we acknowledge that other esolangs may not fall into any of these categories, but this should suffice for a ‘first cut’ at classifying the field.
One key observation is that almost all esolangs have low-level semantics, in terms of primitive arithmetic operations on integer values, limited input/output facilities and primitive control flow. Higher-level facilities such as abstract data types, modularity and concurrency features almost entirely absent.
Another general observation is that the surface-level, concrete syntax of all esolang programs is highly unconventional. Typical esolang code does not look like mainstream code, indeed, it may not look like code at all (Cox, 2013).
@@ 2.1. INTERCAL
INTERCAL is an unusual acronym: confusingly it stands for ‘Compiler Language With No Pronounceable Acronym’. The language was invented in 1972 by two Princeton students as a parody of the verbose imperative languages of the day.
The language reference manual (Woods and Lyon, 1973) is probably the first mention of ‘esoteric’ in relation to programming languages of this nature:
@@
One of the (many) highly original features of INTERCAL is the PLEASE command modifier. The compiler requires occasional politeness, however excessive PLEASEs cause an compile error. In a remarkable back-to-the-future moment, recent ChatGPT users have demonstrated politeness with AI prompts (Yuan et al., 2024).
@@
The C language version of the INTERCAL compiler improved its reach. Now, INTERCAL is available on many open-source package management systems. For instance, you can brew install intercal on macOS.
@@
2.2. Natural Language Languages
Programming in natural language was an early ambition for language developers. Early languages like COBOL reflect this drive, to some extent. Perhaps the most ‘natural’ programming language today is Inform7, used for interactive text adventure games (Nelson, 2006).
@@ 2.2. Natural Language Languages
For such languages, although the syntax is natural, the structured nature of the artifact means that programs can generally be parsed in a straightforward manner. The underlying computational model is generally reasonable, involving integers, simple arithmetic operations, assignments to local variables, conditional tests and basic imperative control flow.
For example, Figure 2 shows the step-by-step execution of a Shakespeare fibonacci calculation. Characters are individual variables storing integer values. In Shakespeare, only two actors can be on the stage at once. Their dialogue involves assigning values to each other. The most complex part of Shakespeare is the representation of integer literal constants, involving complimentary and derogatory nouns (denoting +1 and -1 respectively) with every prefixed adjective doubling a value. So ‘a sweet red rose’ would be $2\times 2\times 1$, whereas a ‘poisonous toad’ would be 2×12$\times$-12 × - 1.
@@ 2.3. Quirky Syntax
Both these languages feature relatively straighforward abstract machines. They are stack based with a sensible instruction set that has appropriate low-level operations. In each case, the only truly unusual aspect of the language is the syntax.
@@
2.4. Quirky Computational Model
@@ 2.5. Discussion
54. Beware of the Turing tar-pit in which everything is possible but nothing of interest is easy.
@@
Cox (Cox, 2013) considers various aspects of esolangs, presenting a philosophical and largely non-technical appreciation. We review further esolang literature in Section 8.
@@ 3. Appeal of Esoteric Languages
Esolangs have an enduring sense of appeal to computer scientists. In this section, we explore the reasons for this popularity. The characteristics we analyse have more to do with the nature of the programmer, than of the programming language. Although it’s natural for one to reflect the other…
@@ 3.1. Playfulness
Really, esolangs are only appropriate for expressing ‘toy’ programs and simple algorithm implementations. For the most part, esolangs feature primitive IO, perhaps only supporting putchar and getchar with no higher-level libraries. As such, code written in esolangs is simple in terms of I/O, probably the kind of program required in coding competitions like the Code Olympics and the Advent of Code. This reinforces the sense of playfulness; many people attempt Advent of Code problems in esolangs. For instance, a google search for ‘advent of code in brainf***’ highlights a number of attempts, demonstrating a non-trivial intersection between people who like esolangs and people who enter coding competitions.
@@
We acknowledge that it is not a new educational idea to connect play and learning. For instance, the Logo turtle graphics language is a playful way for school-age learners to encounter geometry, maths and programming. Papert (Papert, 2002) introduces the notion of ‘hard fun’, which captures ‘hard challenging things to do, and complex artifacts to play with’. Again, this challenge is evident in esolangs.
@@ 3.2. Nostalgia
Again, this attitude harmonizes neatly with the legendary story of Mel the Hacker from Royal McBee, recorded in Eric Raymond’s jargon file (Nather, 1983), describing how Mel wrote optimal code that worked perfectly due to the increment of a self-modifying instruction resulting in a jump at precisely the right moment in the program.
@@
The key takeaway here is the hacker instinct that programming ought to be difficult, and esolangs enforce this. By contrast, coding in mainstream languages is becoming more and more simple. Another reason why coding in mainstream languages is relatively straightforward This has much to do with the tooling support. Many new developers use integrated development environments like Visual Studio Code, featuring syntax highlighting, autocorrection, large language model (LLM) hints, static checking etc. Such modern IDEs have minimal support for esolangs: LOLCODE on VS Code is a notable exception, featuring a syntax highlighting plugin—see Figure 4. for instance. Mostly just text editor and terminal compilation.
@@
Further, many mainstream managed language runtimes feature dynamic safety checks, e.g. for out-of-bounds memory accesses. Even traditionally unsafe languages, like C, can be executed with address sanitisers or binary instrumentation to detect bugs. Esolang runtimes do not have these safety checks, by default.
Another issue relates to AI-generated code. LLMs are notoriously helpful at generating reasonable code for mainstream languages. However AI synthesised code is less successful for esolangs, see Section 7.
@@ 3.3. Sense of Belonging
To the novice programmer, unfamiliar syntax and abstract concepts can feel confusing and discouraging. In contrast, experienced programmers often embrace this complexity, finding enjoyment and stimulation in the challenge. What may seem obscure or unconventional to a beginner can appear refreshingly creative and original to a seasoned practitioner (Cox, 2013). Esoteric programming languages, in particular, foster a sense of community—not necessarily through fluency in a single language, but through a shared appreciation for playful, cryptic syntax and unconventional algorithmic expression. This shared enthusiasm hints at a kind of cultic initiation, where an ‘inner circle’ of programmers engages with code not merely as a practical tool, but as a source of amusement, aesthetic pleasure, and intellectual play.
@@ 3.4. Artistic Expression
We might also mention ‘live programming’, where the programming itself is generating art (e.g. live music (Blackwell and Collins, 2005)). This is usually performative art, rather than artifact creation.
@@ 4. Pedagogic Value
Many esolangs pare away the conveniences of high-level languages and expose directly the underlying computational model. The abstract machine may be some variant of a Turing machine, or some alternative. The programmer is forced to consider how control flow, program state, and logic work at a low level. This understanding is incredibly valuable, as it concretizes what’s happening when programs execute.
@@
Finally, for programming educators, esolangs offer the opportunity to experience similar levels of frustration as the average novice programmer does with a mainstream language. You might be able to code quicksort in a handful of lines of Haskell, but what happens when you try to write a sorting algorithm in an esolang? We might be able to enhance our empathy as educators, as we relive the agony of debugging programs we hardly understand.
@@ 5.1.1. Mostly Students?
Many esolangs were designed by Computer Scientists during their university studies. For instance, the INTERCAL language was created by two Princeton University students: Woods and Lyon. The Whitespace language was created by two undergraduates at Durham University: Brady and Morris. Further, the Shakespeare language was created by students Åslund and Wiberg as a coursework project submission in their Compilers course at Royal Institute of Technology in Stockholm.666 https://web.archive.org/web/20220721085340/http://shakespearelang.sourceforge.net/report/shakespeare/shakespeare.html
@@ 5.1.2. Mostly Males?
Similarly to common perceptions regarding mainstream languages, it seems that the majority of esolang designers are males. This is based on a cursory analysis of designer names, with some implicit cultural assumptions. A more formal analysis would be required to verify this claim. However it generally concurs with the work of Hermans on feminism in programming languages (Hermans and Schlesinger, 2024), as she notes that ‘exclusion of women from technology is both a cause and a result of a gendered interpretation of what programming is and what contributions matter’ Hermans notes that valuing ‘hard’ and employing ‘difficult tools that not everyone can use’ are both male-dominant characteristics, which again, might align with the prevailing trends of esolangs.

### v2

@@ Abstract.
Esoteric programming languages are challenging to learn, but their unusual features and constraints may serve to improve programming ability. From languages designed to be intentionally obtuse (e.g. INTERCAL) to others targeting artistic expression (e.g. Piet) or exploring the nature of computation (e.g. Fractan), there is rich variety in the realm of esoteric programming languages. This essay examines the counterintuitive appeal of esoteric languages and seeks to analyse reasons for this popularity. We will explore why people are attracted to esoteric languages in terms of (a) program comprehension and construction, as well as (b) language design and implementation. Our assertion is that esoteric languages can improve general PL awareness, at the same time as enabling the esoteric programmer to impress their peers with obscure knowledge. We will also consider pedagogic principles and the use of AI, in relation to esoteric languages. Emerging from the specific discussion, we identify a general set of ‘good’ reasons for designing new programming languages. It may not be possible for anyone to be exhaustive on this topic, and it is certain we have not achieved that goal here. However we believe our most important contribution is to draw more attention to the varied and often implicit motivations involved in programming language design.
programming language design, esolangs, programming humor
copyright: ccdoi: 10.1145/3759429.3762632journalyear: 2025isbn: 979-8-4007-2151-9/25/10conference: Proceedings of the 2025 ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software; October 12–18, 2025; Singapore, Singaporebooktitle: Proceedings of the 2025 ACM SIGPLAN International Symposium on New Ideas, New Paradigms, and Reflections on Programming and Software (Onward! ’25), October 12–18, 2025, Singapore, Singaporeccs: Software and its engineering Language typesccs: Social and professional topics Cultural characteristics
@@ 1. Introduction
Despite their name, esoteric programming languages are surprisingly popular in the general Computer Science community. This essay is based on a presentation we have given several times to audiences of academic Computer Scientists interested in programming language design; on each occasion we have encountered keen advocates of esoteric languages like Whitespace and Brainf***. However we have also detected a sense of shame—esoteric language usage is an unexpectedly common ‘guilty secret’. In this work, we seek to analyse potential reasons for these attitudes to esoteric languages. We will argue that Computer Scientists need to take esoteric languages seriously.
@@
An esoteric programming language (hereafter contracted to esolang, following general convention) is ‘a programming language designed to experiment with weird ideas, to be hard to program in, or as a joke, rather than for practical use’. [^1] https://esolangs.org/wiki/Esoteric_programming_language Temkin (Temkin, 2017) characterizes an esoteric programming language as one that is ‘intentionally ununsable, uncomputable or conceptual’. Another definition [^2] https://esoteric.codes/blog/esolangs-as-an-experiential-practice is that ‘esolangs are a class of languages made for reasons other than practical use’.
@@
The claim of this essay is that esoteric languages might be inherently useful in some significant ways, and so they deserve to be studied in a serious manner. The first question we seek to address is: What is the appeal of esolangs, in general? Our discussion will cover the following four high-level reasons for their popularity:
@@
Taking inspiration from Wing’s conceptual framework of Computational Thinking (Wing, 2006), we will ask: what do programmers actually do? We will explore how this characteristic behaviour is manifested in the medium of esolangs. Whereas Wing’s claim is that practitioners in other disciplines can benefit from a grasp of computing principles, our study of esoteric languages leads us to believe that computer programmers can benefit from a wider appreciation of other disciplines. This is Gordon’s argument in his Onward! essay from last year (Gordon, 2024) which takes the complementary stance to Wing, encouraging computer programmers to appreciate natural language and linguistics.
In considering pedagogic principles, we will ask Why is it beneficial to expose learners to esoteric languages? We will explore this question through the lens of variation theory, inter alia.
Although esolangs may seem trivial, they deserve our attention since they can change the way we think about programming and language design. As Perlis (Perlis, 1982) puts it:
Epigram 19: A language that doesn’t affect the way you think about programming, is not worth knowing.
@@ 2. Background
In this section, we briefly survey the field of esolangs. We commence with the prototypical esolang, INTERCAL, in Section 2.1. After this, for the sake of convenience we cluster the languages into broadly related sets, considering esolangs that have a natural language syntax (Section 2.2), those with non-textual syntax (Section 2.3) and those with an exotic computational model (Section 2.4). We acknowledge that these characteristic properties are not mutually exclusive. Further, we acknowledge that other esolangs may not fall into any of these categories, but we hope this is sufficient for an imprecise ‘first cut’ at surveying the field.
One key observation is that almost all esolangs have low-level semantics, in terms of primitive arithmetic operations on integer values, limited input/output facilities and primitive control flow. Higher-level facilities such as abstract data types, modularity and concurrency features are almost entirely absent.
Another general observation is that the surface-level, concrete syntax of all esolang programs is highly unconventional. Typical esolang code does not look like mainstream code, indeed, it may not look like code at all (Cox and McLean, 2012).
@@ 2.1. INTERCAL
We first discuss INTERCAL, since this is the archetype of esolangs. To the best of our knowledge, its language reference manual (Woods and Lyon, 1973) makes the first mention of ‘esoteric’ in relation to programming languages of this nature.
@@
Like many subsequent esolangs, INTERCAL was devised by university students as a playful, humorous artifact. It has become ‘the longest-running and most convoluted joke in the history of programming language design’ (Raymond, 2010). Although the language initially had a minimal compiler and negligible source code corpus, it caught the attention of a small group of dedicated enthusiasts including high profile Computer Scientists like Raymond (Raymond, 1990) and Knuth (Knuth, 2011) who helped to popularize it.
INTERCAL is not a well-formed acronym or abbreviation: confusingly [^3] The contraction is ‘for obvious reasons’, according to the language manual (Woods and Lyon, 1973). it stands for ‘Compiler Language With No Pronounceable Acronym’. The language was invented in 1972 by two Princeton students as a parody of the verbose imperative languages of the day.
INTERCAL has unusual lexical elements, such as the use of the characters ¢ and \$ as operators. There are also unconventional binary operations, such as bitwise interleavings of values. For these reasons among others, fairly standard arithmetic operations become complex and convoluted in INTERCAL.
One of the (many) highly original features of INTERCAL is the PLEASE command modifier. The compiler requires occasional politeness, however excessive PLEASEs cause a compilation error. In a remarkable back-to-the-future moment, recent ChatGPT users have demonstrated similar politeness with AI prompts (Yuan et al., 2024).
@@
The C language version of the INTERCAL compiler improved its reach (Raymond, 1990). Now, INTERCAL is available on many open-source package management systems. For instance, you can brew install intercal on macOS or alternatively, apt install intercal on Ubuntu.
We consider INTERCAL to be an archetypal esolang since it demonstrates foundational principles that define the genre, i.e. syntactic absurdity and a satirical stance toward conventional programming language design. Later esolangs share these playful characteristics. INTERCAL sparked amusement and curiosity among the broader software developer community, which continues to engage with esolangs more generally.
2.2. Natural Language-like Programming Languages
Programming in natural language was an early ambition for language developers. Initial high-level languages like COBOL reflect this drive, to some extent. Perhaps the most ‘natural’ programming language today is Inform7, used for interactive text adventure games (Nelson, 2006).
@@ 2.2. Natural Language-like Programming Languages
For such languages, although the syntax is natural, the fixed structure of the artifact means that programs can be parsed in a straightforward manner. The underlying computational model is generally reasonable, involving integers, simple arithmetic operations, assignments to local variables, conditional tests and basic imperative control flow.
For example, Figure 2 shows the step-by-step execution of a Shakespeare fibonacci calculation. Characters are individual variables storing integer values. In Shakespeare, only two actors can be on the stage at once. Their dialogue involves assigning values to each other. The most complex part of Shakespeare is the representation of integer literal constants, involving complimentary and derogatory nouns (denoting +1 and -1 respectively) with every prefixed adjective doubling a value. So ‘a sweet red rose’ would be $2\times 2\times 1$, whereas a ‘poisonous toad’ would be $2\times-1$.
@@ 2.3. Non-Textual Syntax
Both these languages feature relatively straightforward abstract machines. They are stack based with a sensible instruction set that has appropriate low-level operations. In each case, the only truly unusual aspect of the language is the syntax.
@@
2.4. Unconventional Computational Model
@@ 2.5. Discussion
Epigram 54: Beware of the Turing tar-pit in which everything is possible but nothing of interest is easy.
@@
Cox (Cox and McLean, 2012) considers various aspects of esolangs, presenting a philosophical and largely non-technical appreciation. We review further esolang literature in Section 8.
@@ 3. Appeal of Esoteric Languages
Esolangs have an enduring sense of appeal to many computer scientists, particularly programming language designers. In this section, we explore the reasons for this popularity. The characteristics we analyse have more to do with the nature of the programmer, than of the programming language. Although it’s natural for one to reflect the other …
@@ 3.1. Playfulness
One way to define ‘play’ is an activity which is defined by the activity alone, and is performed to discover the outcome, rather than to achieve a pre-defined goal. In this sense, the playfulness of esolangs is significant.
Really, esolangs are only appropriate for expressing ‘toy’ programs and simple algorithm implementations. For the most part, esolangs feature primitive IO, perhaps only supporting putchar and getchar functions with no higher-level libraries. As such, code written in esolangs is simple in terms of I/O, probably the kind of program required in coding competitions like the Code Olympics and the Advent of Code. This reinforces the sense of playfulness; many people attempt Advent of Code problems in esolangs. For instance, a google search for ‘advent of code in brainf***’ highlights a number of attempts, demonstrating a non-trivial intersection between people who like esolangs and people who enter coding competitions.
@@
We acknowledge that it is not a new educational idea to connect play and learning. For instance, the Logo turtle graphics language is a playful way for school-age learners to encounter geometry, maths and programming. Papert (Papert, 2002) introduces the notion of ‘hard fun’, which captures ‘hard challenging things to do, and complex artifacts to play with’. Further, play is experiential learning by discovery (Kolb, 2014). To players, the activity is key and the outcome is not necessarily important. Again, these factors are evident in esolangs.
Esolangs preserve a ‘sense of fun’ in programming. Perlis, as quoted by Abelson and Sussman (Abelson and Sussman, 1996), states:
I think that it’s extraordinarily important that we in computer science keep fun in computing. When it started out, it was an awful lot of fun …I hope the field of computer science never loses its sense of fun.
@@ 3.2. Nostalgia
Again, this attitude harmonizes neatly with the legendary story of Mel the Hacker from Royal McBee, recorded in Eric Raymond’s jargon file (Nather, 1983), describing how Mel wrote optimal code that worked perfectly on a specific rotating memory machine due to the increment of a self-modifying instruction resulting in a jump at precisely the right moment in the program.
@@
The key takeaway here is the hacker instinct that programming ought to be difficult, and esolangs enforce this. By contrast, coding in mainstream languages is becoming progressively simpler. Tooling support is another reason why coding in mainstream languages is relatively straightforward. Many new developers use integrated development environments like Visual Studio Code, featuring syntax highlighting, autocorrection, large language model (LLM) hints, static checking etc. Such modern IDEs have minimal support for esolangs: LOLCODE on VS Code is a notable exception, featuring a syntax highlighting plugin—see Figure 4 for instance. Mostly developers working with esolangs just use a text editor and terminal compilation.
@@
Further, many mainstream managed language runtimes feature dynamic safety checks, e.g. for out-of-bounds memory accesses. Even traditionally unsafe languages, like C, can be executed with address sanitizers or binary instrumentation to detect bugs. Esolang runtimes do not have these safety checks, by default.
Another issue relates to AI-generated code. LLMs are notoriously helpful at generating reasonable code for mainstream languages. However AI synthesized code is less successful for esolangs, see Section 7.
@@ 3.3. Sense of Belonging
To the novice programmer, unfamiliar syntax and abstract concepts can feel confusing and discouraging. In contrast, experienced programmers often embrace this complexity, finding enjoyment and stimulation in the challenge. What may seem obscure or unconventional to a beginner can appear refreshingly creative and original to a seasoned practitioner (Cox and McLean, 2012). Esoteric programming languages, in particular, foster a sense of community—not necessarily through fluency in a single language, but through a shared appreciation for playful, cryptic syntax and unconventional algorithmic expression. This shared enthusiasm hints at a kind of cultic initiation, where an ‘inner circle’ of programmers engages with code not merely as a practical tool, but as a source of amusement, aesthetic pleasure, and intellectual play.
@@ 3.4. Artistic Expression
We might also mention ‘live coding’, where the programming itself is generating art (e.g. live music (Blackwell and Collins, 2005)). This is usually performative art, rather than artifact creation. More generally, usable live programming involves fluid code editing with immediate feedback (McDirmid, 2013).
@@ 4. Pedagogic Value
Many esolangs pare away the conveniences of high-level languages and expose directly the underlying computational model. The abstract machine may be some variant of a Turing machine, or some alternative. The programmer is forced to consider how control flow, program state, and logic work at a low level. This understanding is highly valuable, as it concretizes what’s happening when programs execute.
@@
Finally, for programming educators, esolangs offer the opportunity to experience similar levels of frustration as the average novice programmer does with a mainstream language. You might be adept at expressing quicksort in a handful of lines of Haskell, but what happens when you try to write a sorting algorithm in an esolang? We might be able to enhance our empathy as educators, as we relive the agony of debugging cryptic fragments of programs we hardly understand.
@@ 5.1.1. Mostly Students?
Many esolangs were designed by Computer Scientists during their university studies. For instance, the INTERCAL language was created by two Princeton University students: Woods and Lyon. The Whitespace language was created by two undergraduates at Durham University: Brady and Morris. Further, the Shakespeare language was created by students Åslund and Wiberg as a coursework project submission in their Compilers course at Royal Institute of Technology in Stockholm.777 https://web.archive.org/web/20220721085340/http://shakespearelang.sourceforge.net/report/shakespeare/shakespeare.html
@@ 5.1.2. Mostly Males?
Similarly to common perceptions regarding mainstream languages, it seems that the majority of esolang designers are males. This is based on a cursory analysis of designer names, with some implicit cultural assumptions. A more formal analysis would be required to verify this claim. However it generally concurs with the work of Hermans on feminism in programming languages (Hermans and Schlesinger, 2024), as she notes that ‘exclusion of women from technology is both a cause and a result of a gendered interpretation of what programming is and what contributions matter’. Hermans notes that valuing ‘hard’ and employing ‘difficult tools that not everyone can use’ are both male-dominant characteristics, which again, might align with the prevailing trends of esolangs.